# Socialized Learning: A Survey of the Paradigm Shift for Edge Intelligence in Networked Systems

### v1

@@ Abstract
Amidst the robust impetus from artificial intelligence (AI) and big data, edge intelligence (EI) has emerged as a nascent computing paradigm, synthesizing AI with edge computing (EC) to become an exemplary solution for unleashing the full potential of AI services. Nonetheless, challenges in communication costs, resource allocation, privacy, and security continue to constrain its proficiency in supporting services with diverse requirements. In response to these issues, this paper introduces socialized learning (SL) as a promising solution, further propelling the advancement of EI. SL is a learning paradigm predicated on social principles and behaviors, aimed at amplifying the collaborative capacity and collective intelligence of agents within the EI system. SL not only enhances the system’s adaptability but also optimizes communication, and networking processes, essential for distributed intelligence across diverse devices and platforms. Therefore, a combination of SL and EI may greatly facilitate the development of collaborative intelligence in the future network. This paper presents the findings of a literature review on the integration of EI and SL, summarizing the latest achievements in existing research on EI and SL. Subsequently, we delve comprehensively into the limitations of EI and how it could benefit from SL. Special emphasis is placed on the communication challenges and networking strategies and other aspects within these systems, underlining the role of optimized network solutions in improving system efficacy. Based on these discussions, we elaborate in detail on three integrated components: socialized architecture, socialized training, and socialized inference, analyzing their strengths and weaknesses. Finally, we identify some possible future applications of combining SL and EI, discuss open problems and suggest some future research, with the hope of arousing the research community’s interest in this emerging and exciting interdisciplinary field.
@@ I-A From Cloud Computing to Edge Intelligence (EI)
Traditional cloud computing sends data from devices like sensors and smartphones to the cloud for processing, then returns results to users. Although this approach can provide abundant computational and storage resources, it also includes difficulties such as latency in the network, bandwidth usage, and privacy and security issues. In response to these issues, a novel computing paradigm, edge computing (EC), has garnered extensive attention in recent years [4]. EC shifts computational services to the network edge, reducing data transmission and enhancing processing efficiency and user experience. The transition from centralized cloud computing to EC underscores the necessity of robust communication frameworks, network design, and resource allocation, catering to the unique demands of edge environments for efficient system operation and data transmission.
@@
Building upon the foundation of EC, EI has been introduced as an emergent paradigm, integrating AI with EC to facilitate intelligent data processing and decision-making at the network edge [5, 6, 7]. EI, employing AI algorithms and models on edge devices, can rapidly perceive and train on vast troves of local data [8]. This enables rapid adaptation to evolving environments, reduced latency, and enhanced computing efficiency. Consequently, EI is envisaged as a formidable driving force in the future development of networks. EI’s reliance on AI algorithms at the network edge also amplifies the need for efficient communication, computing, resource allocation, and other aspects to manage complex data flows and interactions among various devices.
@@ I-B Combination of EI and Socialized Learning (SL)
Despite the immense potential and prospects of EI, its evolution still confronts numerous challenges and dilemmas [11]. For instance, users inevitably generate analogous machine learning (ML) tasks, potentially necessitating identical categories of data or even anticipating congruent training outcomes. This proliferation of repetitive model training precipitates a substantial redundancy in computations across the network and a squandering of finite edge resources. Moreover, smaller data sizes can lead to model overfitting. Consequently, prevalent meaningless model training results in diminished training efficiency and a degradation in service quality.
Another challenge encompasses effectuating efficient communication, resource allocation, and privacy protection within EI. Given the involvement of various types and devices, pivotal issues that demand resolution include coordinating collaboration and competition among them, balancing mutual interests and costs, and ensuring security in interactions [12].
@@
Hence, SL is a learning paradigm based on social principles and behaviors, aimed at enhancing the collaborative capabilities and collective intelligence of various agents within the EI system. SL draws inspiration from a myriad of social interaction modalities present in human societies, such as cooperation, submissive and conflict interactions, and applies them within the EI system. SL perceives each device as an independent, autonomous entity with social awareness. These devices can learn not only from their own data and experiences but also from the data and experiences of others. Through this method of social interaction, SL can effectively utilize data resources within the EI, reduce redundant computations, and enhance training efficiency and service quality. Moreover, SL can effectively allocate computational resources within the EI system, optimize communication costs, and safeguard privacy and security. SL promises to endow the EI system with greater intelligence, adaptability, and robustness.
@@
A real-life application of SL in EI. In the context of intelligent transportation systems, the integration of EI and SL enables autonomous vehicles to learn collaboratively and adapt to dynamic road conditions. A network of self-driving vehicles, equipped with EI capabilities and SL algorithms, continuously gathers and processes real-time data about traffic, road conditions, and weather using onboard sensors and EC resources. This decentralized approach allows for swift responses to immediate situations [14]. The true potential of this system lies in the application of SL principles, enabling vehicles to learn from their own experiences and the collective wisdom of the entire fleet. When a vehicle encounters a novel situation, such as an unexpected obstacle or congestion, it shares this information with other vehicles in the network [15]. This shared knowledge is incorporated into the collective intelligence, benefiting all vehicles.
@@
For instance, if a vehicle detects a hazard, it can alert other vehicles to take alternate routes or proceed with caution. This real-time information sharing contributes to the long-term learning and optimization of the entire transportation network. As more vehicles encounter and share similar situations, the system develops a comprehensive understanding of the city’s traffic dynamics and potential hazards [16]. This collaborative learning approach allows vehicles to proactively adapt to changing conditions based on insights from the fleet. Over time, this learning process enhances the safety, efficiency, and resilience of the autonomous transportation system, leading to smoother traffic flow, reduced congestion, and minimized accidents. The combination of EI and SL exemplifies the transformative potential of these technologies in revolutionizing transportation networks. By leveraging EC and collaborative learning, vehicles make intelligent decisions in real time while contributing to the collective intelligence of the system, ultimately improving urban mobility.
@@ I-C Scenarios Overview of SL in EI
The complex synergy of EI and SL has been integrated and reflected in various fields. This fusion leverages the strengths of both EI and SL, enabling systems to not only process data at the edge with increased efficiency and enhance system efficiency and intelligence. Additionally, it allows systems to learn and adapt through shared experiences and social principles, strengthening their adaptability to complex environments and user experiences.
@@
Autonomous Driving. The integration of EI and SL primarily manifests in the use of edge devices for real-time data processing and decision-making. The deployment on vehicles enables collaborative learning among vehicles, sharing road condition information, thereby enhancing overall driving safety and efficiency. For instance, vehicles can react based on their own sensory inputs and learn from the experiences of other vehicles through communication and data sharing.
IoT. The integration of EI and SL enables devices to process data and share knowledge and experiences through SL algorithms. In smart home systems, this facilitates more accurate environmental control and user behavior prediction, elevating the intelligence of the entire IoT system through device interconnectivity and data sharing.
Smart Healthcare. The combination of EI and SL improves efficiency and accuracy in medical data processing. By deploying edge resources for preliminary data processing and analysis, the burden on central servers is alleviated, while patient data privacy is protected. Additionally, the establishment of edge-assisted medical systems allows for cost savings and improved healthcare service delivery, facilitated by shared knowledge and experiences among medical devices.
@@
Smart Transport and UAVs. EI and SL integration is evident in the use of UAVs for data collection and processing. Deploying EC and SL algorithms on UAVs facilitates real-time monitoring and data processing, improving system response time and efficiency. UAVs sharing flight experiences and data processing strategies through SL enhance the performance of the entire system.
@@ I-D Our Contributions
This paper’s structure, depicted in Fig. 2, commences with an introduction to the backgrounds of EI and SL, followed by a review of pertinent concepts in Sec. II. The rationale for merging EI and SL is in Sec. III, followed by an exposition on the integration points: socialized architecture, training, and inference in Secs. IVVI. Tutorials of SL in EI are presented in Sec. VII. Applications, open problems, and future research are discussed in Sec. VIII, concluding with Sec. IX.
@@ II-A1 Definition of EI
  • AI: AI simulates human intelligence, encompassing techniques like deep learning. The main objective of AI is to model intelligent human behavior by discerning patterns and insights from extensive data repositories.
@@
  • Interplay of AI and EC: EI emerges naturally from the synergy of AI and EC. AI lends advanced techniques to EC, while the latter offers AI-varied application platforms [8]. EC’s processing at the data source reduces latency, critical for analysis of large datasets like audio or video. These rapid insights power applications like public transport planning [19], urban monitoring [20] and early forest fire alerts.
@@ II-C SL: A Function incorporating Multiple Intelligence
  • If we want to implement edge-device cooperative inference in EI, we define a two-layer structure with edge servers and devices. $\mathbf{R}_{a}$ is set to assistance, signifying a dynamic where one layer offers or seeks help from another. Inference operations are denoted as $\mathbf{I}$, and tasks as $\mathbf{L}$.
With this function form, we can flexibly express and design various SL schemes and match them with EI. A visual summary of these schemes and their corresponding parameter values will be presented in the taxonomy graph in end of Section VI (Fig. 37). This flexibility also extends to the optimization of network architectures and computational resource allocation, ensuring that the integration of SL and EI is not only intelligent but also systemically efficient and communication-coherent.
@@ II-D Opportunities Brought by the Integration of EI and SL
SL deepens EI’s understanding of cooperation and improves EI systems’ intelligence and collaboration through human social interaction inspiration. This understanding fosters improved decision-making, addressing gaps that traditional AI might overlook. In decentralized EI, SL harmonizes entities to optimize resources for collective goals. Furthermore, SL’s emphasis on consensus and collective intelligence yields adaptable, resilient systems. This boosts trust and ensures efficient configurations in EI systems.
Conversely, EI’s resources and intelligence offer fertile ground for SL. EI’s distributed architecture enables seamless SL integration. SL entities leverage edge resources to optimize computation and resource allocation. SL models in edge contexts access EI capabilities, creating unique, society-inspired services. Lastly, integrating EI with SL enhances compatibility, scalability, and security, boosting performance. The integration further facilitates a more dynamic resource allocation and network management, enabling the EI system to adapt and respond to varying computational demands and communication needs, thereby exemplifying a comprehensive system-level approach to modern AI and EC challenges.
However, many studies focus on using SL to enhance EI efficiency, neglecting the challenges of SL and the fast advancements in EI methods. This survey emphasizes the synergistic potential of EI and SL, aiming to construct a potent platform for maximizing the value of information, stimulating resource and intelligence sharing at the edge, and enhancing system efficiency.
@@ II-E Review of Existing Similar Concepts
Multi-agent Systems (MAS), Collective Intelligence (CI) and ML Ensembles: Originating from distributed AI [43], the concept of MAS encompasses autonomous agents collaborating, an idea that is further expanded in SL to incorporate the dynamics of society. CI, popularized by Lévy [44], focuses on collaborative efforts for shared objectives, mirroring SL’s emphasis on social collaboration. Introduced by Breiman [45], these methods in ML, such as random forests and boosting algorithms, parallel SL’s approach by integrating diverse models for enhanced learning, akin to SL’s focus on integrating diverse social interactions.
@@ III-A1 Communication Cost
Furthermore, disparate data sizes and non-IID distributions complicate communication. Harnessing SL within EI requires addressing these communication complexities [47]. To address these complexities effectively, exploring innovative collaborative methods for efficient network designs and data transmission is essential. Such methods should align with the unique demands of EI environments, aiming to enhance model synchronization and reduce bandwidth constraints.
@@ III-A3 Privacy and Security
The increased exposure from diverse security protocols and edge heterogeneity necessitates enhanced privacy and security measures against multiple vulnerabilities. Data processed and stored at the edge faces increased exposure from diverse security protocols [47]. The distributed nature of EI introduces multiple vulnerabilities. Diverse threats and edge heterogeneity complicate uniform security measures.
@@ III-B2 SL-Driven Resource Allocation
In the intricate landscape of EI, resource management not only involves communication but also encompasses caching, computing, and control. In EI’s intricate landscape, SL aids in managing the diverse demands of computation, radio resources, and task allocation by leveraging collective knowledge to anticipate and optimize resource use. This socially-aware approach, backed by real-time adaptability, empowers EI systems to meet the dynamic requirements of modern tasks and applications efficiently [50].
@@ III-C Lessons Learned
The convergence of EI and SL illuminates synergistic gains while highlighting critical challenges such as agent coordination and privacy concerns in collaborative settings. The symbiotic potential of EI and SL is evident in its capacity to alleviate communication overhead, streamline resource management and maintain privacy. SL’s approach to collaborative learning navigates through EI’s complexities, offering adaptive problem-solving and efficient communication where traditional methods may falter. This cooperative model not only refines information exchange but also underscores the need for robust and adaptable communication networks and computing architectures, essential for dynamic resource allocation and processing in EI systems.
Resource optimization extends to computing and control processes, and this necessitates harnessing SL’s approach to optimize resources, leveraging shared intelligence for more efficient use. SL also fortifies privacy by sharing insights over data, enhancing network intelligence and thwarting data misuse. Beyond providing immediate solutions, the integration of SL into EI aims to create inherently adaptive systems, poised to evolve with emerging technological demands and ensure long-term relevance.

### v2

@@ Abstract
Amidst the robust impetus from artificial intelligence (AI) and big data, edge intelligence (EI) has emerged as a nascent computing paradigm, synthesizing AI with edge computing (EC) to become an exemplary solution for unleashing the full potential of AI services. Nonetheless, challenges in communication costs, resource allocation, privacy, and security continue to constrain its proficiency in supporting services with diverse requirements. In response to these issues, this paper introduces socialized learning (SL) as a promising solution, further propelling the advancement of EI. SL is a learning paradigm predicated on social principles and behaviors, aimed at amplifying the collaborative capacity and collective intelligence of agents within the EI system. SL not only enhances the system’s adaptability but also optimizes communication, and networking processes, essential for distributed intelligence across diverse devices and platforms. Therefore, a combination of SL and EI may greatly facilitate the development of collaborative intelligence in the future network. This paper presents the findings of a literature review on the integration of EI and SL, summarizing the latest achievements in existing research on EI and SL. Subsequently, we delve comprehensively into the limitations of EI and how it could benefit from SL. Special emphasis is placed on the communication challenges and networking strategies and other aspects within these systems, underlining the role of optimized network solutions in improving system efficiency. Based on these discussions, we elaborate in detail on three integrated components: socialized architecture, socialized training, and socialized inference, analyzing their strengths and weaknesses. Finally, we identify some possible future applications of combining SL and EI, discuss open problems and suggest some future research, with the hope of arousing the research community’s interest in this emerging and exciting interdisciplinary field.
@@ I-A From Cloud Computing to Edge Intelligence (EI)
Traditional cloud computing sends data from devices like sensors and smartphones to the cloud for processing, then returns results to users. Although this approach can provide abundant computational and storage resources, it also includes difficulties such as latency in the network, bandwidth usage, and privacy and security issues. In response to these issues, a novel computing paradigm, edge computing (EC), has acquired extensive attention in recent years [4]. EC shifts computational services to the network edge, reducing data transmission and enhancing processing efficiency and user experience. The transition from centralized cloud computing to EC underscores the necessity of robust communication frameworks, network design, and resource allocation, catering to the unique demands of edge environments for efficient system operation and data transmission.
@@
Building upon the foundation of EC, EI has been introduced as an emergent paradigm, integrating AI with EC to facilitate intelligent data processing and decision-making at the network edge [5, 6, 7]. EI, employing AI algorithms and models on edge devices, can rapidly perceive and train on vast troves of local data [8]. This enables rapid adaptation to evolving environments, reduced latency, and enhanced computing efficiency. Consequently, EI is regarded as a formidable driving force in the future development of networks. EI’s reliance on AI algorithms at the network edge also amplifies the need for efficient communication, computing, resource allocation, and other aspects to manage complex data flows and interactions among various devices.
@@ I-B Combination of EI and Socialized Learning (SL)
Despite the immense potential and prospects of EI, its evolution still confronts numerous challenges and dilemmas [11]. For instance, users inevitably generate analogous machine learning (ML) tasks, potentially necessitating identical categories of data or even anticipating congruent training outcomes. This massive repetition of model training can lead to a substantial redundancy in computations across the network and a squandering of finite edge resources. Moreover, smaller data sizes can lead to model overfitting. Consequently, prevalent meaningless model training results in diminished training efficiency and a degradation in quality of service (QoS).
Another challenge encompasses achieving efficient communication, resource allocation, and privacy protection within EI. Given the involvement of various types and devices, pivotal issues that demand resolution include coordinating collaboration and competition among them, balancing mutual interests and costs, and ensuring security in interactions [12].
@@
Hence, SL is a learning paradigm based on social principles and behaviors, aimed at enhancing the collaborative capabilities and collective intelligence of various agents within the EI system. SL draws inspiration from a myriad of social interaction modalities present in human societies, such as cooperation, submissive and conflict interactions, and applies them within the EI system. SL perceives each device as an independent, autonomous entity with social awareness. These devices can learn not only from their own data and experiences but also from the data and experiences of others. Through this method of social interaction, SL can effectively utilize data resources within the EI, reduce redundant computations, and enhance training efficiency and QoS. Moreover, SL can effectively allocate computational resources within the EI system, optimize communication costs, and safeguard privacy and security. SL promises to endow the EI system with greater intelligence, adaptability, and robustness.
@@
A real-life application of SL in EI. In the context of intelligent transportation systems, the integration of EI and SL enables autonomous vehicles to learn collaboratively and adapt to dynamic road conditions. A network of self-driving vehicles, equipped with EI capabilities and SL algorithms, continuously gathers and processes real-time data about traffic, road conditions, and weather using onboard sensors and EC resources. This decentralized approach allows for swift responses to immediate situations [14]. The true potential of this system lies in the application of SL principles, enabling vehicles to learn from their own experiences and the collective wisdom of the entire system. When a vehicle encounters a novel situation, such as an unexpected obstacle or congestion, it shares this information with other vehicles in the network [15]. This shared knowledge is incorporated into the collective intelligence, benefiting all vehicles.
@@
For instance, if a vehicle detects a hazard, it can alert other vehicles to take alternate routes or proceed with caution. This real-time information sharing contributes to the long-term learning and optimization of the entire transportation network. As more vehicles encounter and share similar situations, the system develops a comprehensive understanding of the city’s traffic dynamics and potential hazards [16]. This collaborative learning approach allows vehicles to proactively adapt to changing conditions based on insights from the fleet. Over time, this learning process enhances the safety, efficiency, and resilience of the autonomous transportation system, leading to smoother traffic flow, reduced congestion, and minimized accidents. The combination of EI and SL exemplifies the transformative potential of these technologies in revolutionizing transportation networks. By leveraging EC and SL, vehicles make intelligent decisions in real time while contributing to the collective intelligence of the system, ultimately improving urban transportation.
@@ I-C Scenarios Overview of SL in EI
The complex synergy of EI and SL has been integrated and reflected in various fields. This fusion leverages the strengths of both EI and SL, enabling systems to process data at the edge efficiently and enhance system intelligence. Additionally, it allows systems to learn and adapt through shared experiences and social principles, strengthening their adaptability to complex environments and user experiences.
@@
Autonomous Driving. The integration of EI and SL primarily manifests in the use of edge devices for real-time data processing and decision-making. The deployment on vehicles enables collaborative learning among vehicles, sharing road condition information, thereby enhancing overall driving safety and efficiency. Vehicles can react based on their own sensory inputs and learn from other vehicles’ experiences through communication and data sharing.
IoT. The integration of EI and SL enables devices to process data and share knowledge and experiences through SL algorithms. In smart home systems, this facilitates accurate environmental control and user behavior prediction, elevating IoT system intelligence through device interconnectivity and data sharing.
Smart Healthcare. The combination of EI and SL improves efficiency and accuracy in medical data processing. By deploying edge resources for preliminary data processing and analysis, the burden on central servers is alleviated, while patient data privacy is protected. Edge-assisted medical systems allow for cost savings and improved healthcare service delivery, facilitated by shared knowledge and experiences among medical devices.
@@
Smart Transport and Unmanned Aerial Vehicles (UAVs). EI and SL integration is evident in the use of UAVs for data collection and processing. Deploying EC and SL algorithms on UAVs facilitates real-time monitoring and data processing, improving system response time and efficiency. UAVs sharing flight experiences and data processing strategies through SL enhance the performance of the entire system.
@@ I-D Our Contributions
This paper’s structure, depicted in Fig. 2, commences with an introduction to the backgrounds of EI and SL, followed by a review of relevant concepts in Sec. II. The rationale for integrating EI and SL is in Sec. III, followed by an exposition on the integration points: socialized architecture, training, and inference in Secs. IVVI. Tutorials of SL in EI are presented in Sec. VII. Applications, open problems, and future research are discussed in Sec. VIII, concluding with Sec. IX.
@@ II-A1 Definition of EI
  • AI: AI simulates human intelligence, encompassing techniques like deep learning (DL). The main objective of AI is to model intelligent human behavior by discerning patterns and insights from extensive data repositories.
@@
  • Interplay of AI and EC: EI emerges naturally from the synergy of AI and EC. AI provides advanced techniques, while EC supplies varied, real-time application platforms [8]. EC’s processing at the data source reduces latency, which is critical for analyzing large datasets such as audio and video. These rapid insights power applications like public transport planning [19], urban monitoring [20] and early forest fire alerts.
@@ II-C SL: A Function incorporating Multiple Intelligence
  • If we want to implement edge-device collaborative inference in EI, we define a two-layer structure with edge servers (ESs) and devices. $\mathbf{R}_{a}$ is set to assistance, signifying a dynamic where one layer offers or seeks help from another. Inference operations are denoted as $\mathbf{I}$, and tasks as $\mathbf{L}$.
With this function form, we can flexibly express and design various SL schemes and match them with EI. A visual summary of these schemes and their corresponding parameter values will be presented in the taxonomy graph in the end of Section VI, as shown in Fig. 37. This flexibility also extends to the optimization of network architectures and computational resource allocation, ensuring that the integration of SL and EI is not only intelligent but also systemically efficient and communication-coherent.
@@ II-D Opportunities Brought by the Integration of EI and SL
Several recent surveys have investigated the integration of EC with advanced technologies. Wang et al. [47] discussed the convergence of EC and DL, including application scenarios and enabling technologies. Taleb et al. [48] examined 5G network edge cloud architecture. Yang et al. [49] explored the integration of blockchain with EC, addressing key motivations and challenges. Lim et al. [15] reviewed FL in EC, focusing on communication and privacy challenges. Lastly, Ning et al. [20] analyzed EC and ML applications for the Internet of UAVs. While these surveys offer valuable insights, they do not organize and present state-of-the-art research in EI from an SL perspective, providing an innovative and comprehensive understanding of the challenges, techniques, and applications in EI systems.
SL deepens EI’s understanding of cooperation and improves EI systems’ intelligence and collaboration through inspiration from human social interaction. This understanding fosters improved decision-making, addressing gaps that traditional AI might overlook. In decentralized EI, SL harmonizes entities to optimize resources for collective goals. Furthermore, SL’s emphasis on consensus and collective intelligence yields adaptable, resilient systems. This boosts trust and ensures efficient configurations in EI systems.
Conversely, EI’s resources and intelligence offer fertile ground for SL. EI’s distributed architecture enables seamless SL integration. SL entities leverage edge resources to optimize computation and resource allocation. SL models in edge contexts access EI capabilities, creating unique, society-inspired services. Lastly, integrating EI with SL enhances compatibility, scalability, and security, boosting performance.
The integration further facilitates a more dynamic resource allocation and network management, enabling the EI system to adapt and respond to varying computational demands and communication needs, thereby exemplifying a comprehensive system-level approach to modern AI and EC challenges. This survey emphasizes the synergistic potential of EI and SL, aiming to construct a potent platform for maximizing the value of information, stimulating resource and intelligence sharing at the edge, and enhancing system efficiency.
@@ II-E Review of Existing Similar Concepts
Multi-agent Systems (MAS), Collective Intelligence (CI) and ML Ensembles: Originating from distributed AI [43], MAS encompasses autonomous agents collaborating, an idea further expanded in SL to incorporate social dynamics. CI, popularized by Lévy [44], focuses on collaborative efforts for shared objectives, mirroring SL’s emphasis on social collaboration. Introduced by Breiman [45], these methods in ML, such as random forests and boosting algorithms, parallel SL’s approach by integrating diverse models for enhanced learning, akin to SL’s focus on integrating diverse social interactions.
@@ III-A1 Communication Cost
Furthermore, disparate data sizes and non-independent and identically distributed (non-IID) distributions complicate communication. Harnessing SL within EI requires addressing these communication complexities [50]. To address these complexities effectively, exploring innovative collaborative methods for efficient network designs and data transmission is essential. Such methods should align with the unique demands of EI environments, aiming to enhance model synchronization and reduce bandwidth constraints.
@@ III-A3 Privacy and Security
The increased exposure to diverse security protocols and edge heterogeneity necessitates enhanced privacy and security measures against multiple vulnerabilities. Data processed and stored at the edge faces increased exposure from diverse security protocols [50]. The distributed nature of EI introduces multiple vulnerabilities. Diverse threats and edge heterogeneity complicate uniform security measures.
@@ III-B2 SL-driven Resource Allocation
In the intricate landscape of EI, resource management not only involves communication but also encompasses caching, computing, and control. In EI’s intricate landscape, SL aids in managing the diverse demands of computation, radio resources, and task allocation by leveraging collective knowledge to anticipate and optimize resource use. This socialized approach, backed by real-time adaptability, empowers EI systems to meet the dynamic requirements of modern tasks and applications efficiently [53].
@@ III-C Lessons Learned
The convergence of EI and SL illuminates synergistic gains while highlighting critical challenges such as agent coordination and privacy concerns in collaborative settings. The symbiotic potential of EI and SL is evident in its capacity to alleviate communication overhead, streamline resource management and maintain privacy. SL’s approach to collaborative learning navigates through EI’s complexities, offering adaptive problem-solving and efficient communication where traditional methods may falter. This cooperative paradigm not only refines information exchange but also underscores the need for robust and adaptable communication networks and computing architectures, essential for dynamic resource allocation and processing in EI systems.
Resource optimization extends to computing and control processes, and this necessitates harnessing SL’s approach to optimize resources, leveraging shared intelligence to improve usage efficiency. SL also protects privacy by sharing insights over data, enhancing network intelligence and stopping data misuse. Beyond providing immediate solutions, the integration of SL into EI aims to create inherently adaptive systems, poised to evolve with emerging technological demands and ensure long-term relevance.